OpenStudy (anonymous):

ln2 with Taylor series

OpenStudy (phi):

Try expanding the series about 1 \[f(x) = \sum_{k=0}^{\infty} \frac{ f^{(k)}(a) }{k! }(x-a)^k\] i.e. let a=1. \[f(x) = \sum_{k=0}^{\infty} \frac{ f^{(k)}(1) }{k! }(x-1)^k\] if you want f(2), you have \[ f(2) = \sum_{k=0}^{\infty} \frac{ f^{(k)}(1) }{k! }(2-1)^k \\ = \sum_{k=0}^{\infty} \frac{ f^{(k)}(1) }{k! }1^k = \sum_{k=0}^{\infty} \frac{ f^{(k)}(1) }{k! } \] now you need the first and higher order derivatives of ln(x) evaluated at x=a i.e. x=1 f' = 1/x = 1 f'' = -x^(-2) = -1 f'''= +2x^(-3) = +1 \cdot 2! f''''= -3*2 x^(-4)= -1 3! f''''' = +4*3*2 x^(-5) = +1 / 4! or in general \[ f^{(k)}(1) = (-1)^{k+1} \cdot (k-1)!\cdot x^{-k} = (-1)^{k+1} \cdot (k-1)! \]